# Gauge Symmetries, Contact Reduction, and Singular Field Theories

### v1

@@ VII.3 Restriction &\& Reduction
Having shown how we obtain the contact Hamiltonian equations of motion by first contact reducing the pre-symplectic system, and then applying the appropriate constraint algorithm, we shall now consider the reverse process, in which we first carry out the restriction at the symplectic level, before seeking a scaling symmetry with which to contact reduce the constrained theory.
We begin, once again, from the pre-symplectic system $(M_{0},\omega_{0},H_{0})$, and carry out the constraint procedure presented in section (III.1), deducing the final submanifold $M_{f}\hookrightarrow T^{*}Q$. In order to be able to carry out a contact reduction via scaling symmetry, we must either introduce a particular choice of gauge-fixing, or formally quotient out by the action of the gauge transformations; since we are concerned with the dynamics of observable ($i.e$ gauge-invariant) quantities, this choice is somewhat inconsequential, and we denote the physical phase space $\mathcal{P}$, with $\jmath_{\scriptscriptstyle\mathcal{P}}:\mathcal{P}\hookrightarrow M_{f}$.
@@
in which $X_{f}$ refers to the Hamiltonian vector field associated to $f$ via $\iota_{X_{f}}\omega_{\scriptscriptstyle\mathcal{P}}=df$, and similarly for $X_{H_{\scriptscriptstyle\mathcal{P}}}$.
Assuming that the flow of $Z$ acts freely and properly on $\mathcal{P}$, we have a well-defined quotient space $\mathcal{P}/\hskip-2.84526pt\sim$, with submersion $\sigma:\mathcal{P}\rightarrow\mathcal{P}/\hskip-2.84526pt\sim\,$; if $\rho:\mathcal{P}\rightarrow\mathbb{R}_{+}$ is a global scaling function on $\mathcal{P}$, the contact form and Hamiltonian are given by
@@
From the constructions given above, it should be clear that the pre-symplectic and pre-contact constraint algorithms run in parallel. At the $i^{\textrm{th}}$ stage of the pre-symplectic algorithm, the submanifold $M_{i}$, with inclusion $\jmath_{i}:M_{i}\hookrightarrow M_{i-1}$, may be described as the zero-set of the functions $\phi^{\alpha}_{(i)}$. The projection of these $\phi^{\alpha}_{(i)}$ defines a second set of functions $\gamma^{\alpha}_{(i)}$ according to
@@
Figure 1: Commutative diagram showing how the two constraint algorithms run in parallel. Starting from the cotangent bundle, with its canonical symplectic form, the right-hand path shows how we may restrict to the primary constraint manifold $M_{0}$, and carry out the pre-symplectic algorithm, to reach the final space $\mathcal{P}$. This manifold is symplectic, and thus allows a contact reduction to be carried out. Alternatively, the scaling degree of freedom may be excised from the full symplectic phase space, defining the contact manifold $C$. Upon projecting the primary constraint functions that define $M_{0}$, $C$ is restricted to the pre-contact space $C_{0}$, and the left-hand path illustrates the implementation of the pre-contact algorithm. Both paths lead to the same final reduced space $C_{\scriptscriptstyle\mathcal{P}}$.
@@ VIII.1 Restriction &\& Reduction
which is of constant rank 3, provided $u\neq 0$, which we have assumed. This Lagrangian admits a fairly simple scaling symmetry of the form
$\Sigma=\frac{1}{2}\left(u\frac{\partial}{\partial u}+\dot{u}\frac{\partial}{\partial\dot{u}}\right)$ (63)
Making the change of variable $u\mapsto e^{\rho/2}$, the scaling symmetry becomes $\Sigma=\partial_{\rho}$, while the Lagrangian adopts the form
@@
Any vector field solution $X_{H}$ to the geometrical equation $\iota_{X_{H}}\omega_{0}-dH_{0}=0$, restricted to $M_{1}$, must remain tangent to this surface, in order to be considered physically meaningful. As we have seen, this requirement is entirely equivalent to imposing that $\langle(dH_{0})_{p},T_{p}M_{1}^{\perp}\rangle=0$ for all points $p\in M_{1}$. It is straightforward to verify that
@@
Clearly, this is invertible, which indicates that $\psi^{1}$ and $\psi^{2}$ constitute a pair of second class constraints; it is therefore unnecessary to eliminate gauge degrees of freedom, and we henceforth refer to $M_{1}$ as $\mathcal{P}$. The inverse $J^{-1}$ is given by
@@
In order to deduce the dynamical evolution of phase space functions, we should construct the total Hamiltonian; however, having identified $\psi^{1}$ and $\psi^{2}$ as second class, it follows that the total Hamiltonian coincides with $H_{0}$. It is then straightforward to show that the equations of motion for our phase space variables, which follow from $\dot{f}=\{f,H_{0}\}_{D}|_{\scriptscriptstyle\mathcal{P}}$ are
@@ VIII.2 Reduction + Restriction
In this example, we have provided an explicit illustration of how a constrained Hamiltonian system may be reduced to a simpler, dynamically-equivalent theory, by identifying and excising an unphysical scaling degree of freedom. Whilst relatively simple in nature, the system considered has allowed us to present a fully-worked example of the commutativity of contact reduction and phase space restriction, and thus concludes our treatment of particle dynamics. In the following sections, we introduce the geometrical framework required to describe singular field theories, which will then allow us to provide a field-theoretic discussion, which parallels the ideas presented thus far.
@@ IX Multisymplectic Field Theory
The use of multisymplectic geometry to describe classical field theories is an area of active research interest, as the fibred manifolds it employs provide an arena in which a manifestly covariant formalism may be developed, in a finite-dimensional setting [83, 84, 85]. Scaling symmetries of classical field theories have been studied in [34], and a mathematical framework for the contact reduction of multisymplectic theories developed in [29]. The multisymplectic Hamiltonian formalism of singular field theories is still not entirely understood; for example, it is known that there exist certain classes of Lagrangian, for which the construction of the corresponding Hamiltonian is either ambiguous, or simply ill-defined [16]. Additionally, multisymplectic manifolds are not, in general, equipped with a well-defined bracket structure [86], and the local constraint algorithm employs geometrical principles, that render the use of such structures redundant [21, 22]. As a result, in contrast to our study of systems of particles, our analysis of singular field theories will favour the Lagrangian formalism, and in what follows, we shall provide a heavily abridged summary of the most pertinent ideas, referring to [29] for more details.
@@ IX.1 Lagrangian Field Theory
If $\Omega$ is closed but 1-degenerate, we refer to the pair $(\mathcal{M},\Omega)$ as a pre-multisymplectic manifold; such objects will be of particular relevance in our study of singular theories.
In order to formulate the field equations of a dynamical system in a geometrical fashion, we require a class of objects known as multivector fields. In general, a multivector field of degree $r$ on $\mathcal{M}$ is a section $\boldsymbol{X}\in\Gamma(\wedge^{r}T\mathcal{M})$ of the $r^{\textrm{th}}$ exterior power of the tangent bundle. We denote the space of all such multivector fields $\mathfrak{X}^{r}(\mathcal{M}):=\Gamma(\wedge^{r}\,T\mathcal{M})$, and declare some $\boldsymbol{X}\in\mathfrak{X}^{r}(\mathcal{M})$ to be locally decomposable if, for every point $p\in\mathcal{M}$, there exists an open neighbourhood $\mathcal{U}_{p}\subset\mathcal{M}$, and vector fields $X_{1},\,\cdots\,,X_{r}\in\mathfrak{X}^{\infty}(\mathcal{U}_{p})$, such that
@@
Consider a fibre bundle $\pi:E\rightarrow M$ over the $d$-dimensional spacetime manifold $M$; we shall suppose that $M$ is orientable, with volume form $\omega$. Local coordinates on $M$ are written as $(x^{\mu})$, with $\mu=0,\,\cdots,d-1$, so that $\omega=dx^{0}\wedge\,\cdots\wedge dx^{d-1}:=d^{d}x$. The $(n+d)$-dimensional manifold $E$ is referred to as the covariant configuration space, and the first jet bundle $\kappa:J^{1}E\rightarrow E$ of sections of $\pi$ is the natural space upon which to introduce a Lagrangian density [87, 88]; local coordinates on $J^{1}E$, adapted to the bundle structure of $E$, are given by $(x^{\mu},y^{a},y^{a}_{\mu})$, with $a=1,\,\cdots,n$. Introducing the bundle projection $\widehat{\pi}:=\pi\circ\kappa:J^{1}E\rightarrow M$, the Lagrangian density $\mathcal{L}$ may be expressed as a $\widehat{\pi}-$semibasic $d$-form on $J^{1}E$
@@ X Pre-multicontact Systems
In our treatment of scaling symmetries of pre-symplectic systems, we observed that the elimination of a scaling degree of freedom and phase space restriction were commutative processes; indeed, we dedicated significant effort to the construction of the final symmetry-reduced theory via both paths of the commutative diagram 1. When considering field theories, we shall not attempt to retain this same level of generality, and so will restrict our attention to the implementation of a constraint algorithm only after the scaling degree of freedom has been eliminated.
The symmetry reduction of multisymplectic field theories has been studied in detail in [29]; there, we presented a brief account of multicontact geometry, as pertained to the reduction of regular systems. When the theories of interest are singular, far more care must be taken, for it is not guaranteed that the symmetry-reduced form derived from the pre-multisymplectic $\Theta_{L}$ defines a pre-multicontact structure. Consequently, as an additional step, we must ensure that the reduced theory possesses a well-defined Reeb distribution; such considerations will require the introduction of additional mathematical structures, that were not required in [29]. Our discussion of multicontact geometry shall be limited to those concepts which are fundamental to the construction of classical field theories; more details may be found in [26, 25, 27], for example.
@@ X.1 Multicontact Lagrangian Field Theory
Dissipative (or frictional) field theories are the subject of notably less research interest than their conservative counterparts; this is perhaps due to a tendency to wish to embed any such non-conservative configuration within a larger system, often denominated ‘the environment’. The frictional behaviour of the original system is then palliated by the observation that any apparent losses may simply be attributed to an exchange of energy between the system and environment. The philosophical stance which motivates our study of scaling symmetries is diametrically opposed to the notion of creating larger, redundant mathematical structures, with the sole objective of streamlining practical calculations [77, 80, 90]. We posit that the use of the smallest amount of dynamically-irrelevant structure, even at the expense of superficial mathematical simplicity, should be prioritised, wherever possible [91, 92]. In excising superfluous scaling degrees of freedom, seeking a minimal ontology, we are forced to work with theories which are inherently frictional in nature; consequently, we shall dedicate this section to developing the geometrical tools required to describe such systems.
@@
does not necessarily define a pre-multicontact structure. It is therefore necessary to introduce a number of additional concepts, which allow us to determine when a given Lagrangian $d$-form does in fact give rise to such a structure.
We begin by considering a generalisation of the annihilator of a given subspace, that we first encountered in section (II); for the purpose of defining the pre-multicontact structures of this section, in addition to the Lagrangian $d$-form $\Theta_{L}$, we also introduce the notation $\Xi:=\beta^{*}\omega\in\Omega^{d}(\mathcal{S})$; this volume form serves as a reference object, while $\Theta_{L}$ endows the space $\mathcal{S}$ with the formal geometrical structure of a pre-multicontact manifold. Given a regular distribution $\mathcal{D}\subset T\mathcal{S}$, we introduce the space of $r$-forms which annihilate all sections of $\mathcal{D}$
@@
Sections of $\mathcal{D}^{R}$ are the Reeb vector fields; we denote this space $\mathfrak{R}:=\Gamma(\mathcal{D}^{R})$. In practice, we shall always work with coordinates on $\mathcal{S}$ adapted to the bundle structure $E\rightarrow M$, and so $\Xi=dx^{0}\wedge\,\cdots\,\wedge dx^{d-1}$ is a closed form, from which it follows that $\textrm{ker}\,\Xi$ consists of those vector fields whose vertical components $V^{\mu}\partial_{x^{\mu}}$ are vanishing. The space $\textrm{Ann}^{d}(\textrm{ker}\,\Xi)$ then comprises all $d$-forms $\xi\in\Omega^{d}(\mathcal{S})$, for which $\iota_{X}\xi=0$ for any choice of vector field of the form
@@
In addition to the Reeb distribution, we also have the characteristic distribution $\mathcal{C}$, defined as the intersection $\mathcal{C}:=\textrm{ker}\,\Xi\cap\textrm{ker}\,\Theta_{L}\cap\textrm{ker}\,d\Theta_{L}$.
With these geometrical constructions in-hand, the conditions under which the triple $(\mathcal{S},\Theta_{L},\Xi)$ constitutes a pre-multicontact manifold are that, for some $0<k\leqslant n(1+d)$
@@
Note that we have excluded the possibility that $k=0$, for this is precisely the condition required for $(\mathcal{S},\Theta_{L},\Xi)$ to be multicontact, and not pre-multicontact. In practice, when working in adapted coordinates, the local Reeb fields are deduced from $\iota_{R_{\mu}}\Theta_{L}=d^{d-1}x_{\mu}$; however, observe that this does not determine the $R_{\mu}$ uniquely, for it is always possible to add to $R_{\mu}$ an element of the characteristic distribution, that is to say $\mathfrak{R}=\textrm{span}\hskip 1.13809pt(R_{\mu})+\mathcal{C}$.
Given an $r$-form $\xi\in\Omega^{r}(\mathcal{S})$, we introduce the following operator
$\begin{split}\bar{d}\;:\;\Omega^{r}(\mathcal{S})\;&\longrightarrow\;\Omega^{r+1}(\mathcal{S})\\ \xi\;&\longmapsto\;\bar{d}\xi:=d\xi+\sigma_{\Theta_{L}}\wedge\,\xi\end{split}$
where $\sigma_{\Theta_{L}}\in\Omega^{1}(\mathcal{S})$ is known as the dissipation form, and is expressed in local coordinates as
@@
$\Omega_{L}=\bar{d}\Theta_{L}=d\Theta_{L}-\frac{\partial L}{\partial s^{\mu}}\,dx^{\mu}\wedge\,\Theta_{L}$ (92)

### v2

@@ VII.3 Restriction &\& Reduction
In order to carry out the reverse process, in which we seek a scaling symmetry within the constrained theory, we take the pre-symplectic system $(M_{0},\omega_{0},H_{0})$, and implement the procedure presented in section (III.1), deducing the final submanifold $M_{f}\hookrightarrow T^{*}Q$. In order to be able to carry out a contact reduction, we must either introduce a choice of gauge-fixing, or formally quotient out by the action of the gauge transformations; since we are concerned with the dynamics of observable, gauge-invariant quantities, this choice is somewhat inconsequential, and we denote the physical phase space $\mathcal{P}$, with $\jmath_{\scriptscriptstyle\mathcal{P}}:\mathcal{P}\hookrightarrow M_{f}$.
@@
in which $X_{f}$ refers to the Hamiltonian vector field associated to $f$ via $\iota_{X_{f}}\omega_{\scriptscriptstyle\mathcal{P}}=df$, and similarly for $X_{H_{\scriptscriptstyle\mathcal{P}}}$. Assuming that the flow of $Z$ acts freely and properly on $\mathcal{P}$, we have a well-defined quotient space $\mathcal{P}/\hskip-2.84526pt\sim$, with submersion $\sigma:\mathcal{P}\rightarrow\mathcal{P}/\hskip-2.84526pt\sim\,$; if $\rho:\mathcal{P}\rightarrow\mathbb{R}_{+}$ is a global scaling function on $\mathcal{P}$, the contact form and Hamiltonian are given by
@@
From the constructions given above, it should be clear that the pre-symplectic and pre-contact constraint algorithms run in parallel. At the $i^{\textrm{th}}$ stage of the former, the submanifold $M_{i}$, with inclusion $\jmath_{i}:M_{i}\hookrightarrow M_{i-1}$, may be described as the zero-set of the functions $\phi^{\alpha}_{(i)}$, whose projection defines a second set of functions $\gamma^{\alpha}_{(i)}$
@@
Figure 1: Commutative diagram showing how the two constraint algorithms run in parallel. Starting from the cotangent bundle, the right-hand path shows implementation of the pre-symplectic algorithm, to deduce the final space $\mathcal{P}$. This manifold is symplectic, and thus allows a contact reduction to be carried out. Alternatively, the scaling variable may be excised from the full symplectic phase space, defining the contact manifold $C$. Upon projecting the primary constraint functions that define $M_{0}$, $C$ is restricted to the pre-contact space $C_{0}$, and the left-hand path depicts the implementation of the pre-contact algorithm. Both paths lead to the same final reduced space $C_{\scriptscriptstyle\mathcal{P}}$.
@@ VIII.1 Restriction &\& Reduction
which is of constant rank 3, provided $u\neq 0$, which we have assumed. Making the change of variable $u\mapsto e^{\rho/2}$, the Lagrangian adopts the form
@@
Any vector field solution $X_{H}$ to the geometrical equation $\iota_{X_{H}}\omega_{0}=dH_{0}$, restricted to $M_{1}$, must remain tangent to this surface in order to be physically meaningful. As we have seen, this requirement is entirely equivalent to imposing that $\langle(dH_{0})_{p},T_{p}M_{1}^{\perp}\rangle=0$ for all points $p\in M_{1}$. It is straightforward to verify that
@@
Clearly, this is invertible, indicating that $\psi^{1}$ and $\psi^{2}$ constitute a pair of second class constraints; it is therefore unnecessary to eliminate gauge degrees of freedom, and we henceforth refer to $M_{1}$ as $\mathcal{P}$. The inverse $J^{-1}$ is given by
@@
Since $\psi^{1}$ and $\psi^{2}$ are second class constraints, the total Hamiltonian coincides with $H_{0}$. The equations of motion for our phase space variables follow from $\dot{f}=\{f,H_{0}\}_{D}|_{\scriptscriptstyle\mathcal{P}}$, and we find
@@ VIII.2 Reduction + Restriction
In this example, we have provided an explicit illustration of how a constrained Hamiltonian system may be reduced to a simpler, dynamically-equivalent theory, by identifying and excising an unphysical scaling degree of freedom. Whilst relatively simple in nature, the system considered has allowed us to present a fully-worked example of the commutativity of contact reduction and phase space restriction, and thus concludes our treatment of particle dynamics.
@@ IX Multisymplectic Field Theory
The use of multisymplectic geometry to describe classical field theories is an area of active research interest, as the fibred manifolds it employs provide an arena in which a manifestly covariant formalism may be developed, in a finite-dimensional setting [73, 74, 75]. Scaling symmetries of classical field theories have been studied in [34], and a framework for the contact reduction of multisymplectic theories developed in [29]. The multisymplectic Hamiltonian formalism of singular field theories is still not well understood; for example, there exist certain classes of Lagrangian, for which the construction of the corresponding Hamiltonian is either ambiguous, or simply ill-defined [16]. Additionally, multisymplectic manifolds are not, in general, equipped with a well-defined bracket structure [76], and since the local constraint algorithm dispenses with such structures, our analysis of singular field theories will favour the Lagrangian formalism, rather than the Hamiltonian picture employed for particles [21, 22]. In what follows, we provide a heavily abridged summary of the most pertinent ideas, referring to [29] for further details.
@@ IX.1 Lagrangian Field Theory
If $\Omega$ is closed but 1-degenerate, we refer to the pair $(\mathcal{M},\Omega)$ as a pre-multisymplectic manifold. The field equations of a dynamical system are expressed geometrically in terms of multivector fields. In general, a multivector field of degree $r$ on $\mathcal{M}$ is a section $\boldsymbol{X}\in\Gamma(\wedge^{r}T\mathcal{M})$ of the $r^{\textrm{th}}$ exterior power of the tangent bundle. We denote the space of all such multivector fields $\mathfrak{X}^{r}(\mathcal{M}):=\Gamma(\wedge^{r}\,T\mathcal{M})$, and declare some $\boldsymbol{X}\in\mathfrak{X}^{r}(\mathcal{M})$ to be locally decomposable if, around point $p\in\mathcal{M}$, there exists an open neighbourhood $\mathcal{U}_{p}\subset\mathcal{M}$, and vector fields $X_{1},\,\cdots\,,X_{r}\in\mathfrak{X}^{\infty}(\mathcal{U}_{p})$, such that
@@
Consider a fibre bundle $\pi:E\rightarrow M$ over the $d$-dimensional spacetime manifold $M$; we shall suppose that $M$ is orientable, with volume form $\omega$. Local coordinates on $M$ are denoted $(x^{\mu})$, with $\mu=0,\,\cdots,d-1$, so that $\omega=dx^{0}\wedge\,\cdots\wedge dx^{d-1}:=d^{d}x$. The $(n+d)$-dimensional manifold $E$ is referred to as the covariant configuration space, and the first jet bundle $\kappa:J^{1}E\rightarrow E$ of sections of $\pi$ is the natural space upon which to introduce a Lagrangian density [77, 78]; local adapted coordinates on $J^{1}E$ are given by $(x^{\mu},y^{a},y^{a}_{\mu})$, with $a=1,\,\cdots,n$. Introducing the bundle projection $\widehat{\pi}:=\pi\circ\kappa:J^{1}E\rightarrow M$, the Lagrangian density $\mathcal{L}$ may be expressed as a $\widehat{\pi}-$semibasic $d$-form on $J^{1}E$
@@ X Pre-multicontact Systems
In our treatment of pre-symplectic systems, significant effort was dedicated to studying the commutative relationship between the elimination of a scaling degree of freedom and phase space restriction. When considering field theories, we shall not attempt to retain this level of generality, and so will restrict our attention to the implementation of a constraint algorithm only after the scaling degree of freedom has been eliminated.
@@ X.1 Multicontact Lagrangian Field Theory
Dissipative (or frictional) field theories are the subject of notably less research interest than their conservative counterparts; this is perhaps due to a tendency to wish to embed any such non-conservative configuration within a larger system, often denominated ‘the environment’. The frictional behaviour of the original system is then palliated by the observation that any apparent losses may simply be attributed to an exchange of energy between the system and environment. The philosophical stance which motivates our study of scaling symmetries is diametrically opposed to the creation of larger redundant mathematical structures, with the sole objective of streamlining calculations [67, 70, 80]. We posit that the use of the smallest amount of dynamically-irrelevant structure, even at the expense of superficial mathematical simplicity, should be prioritised, where possible [81, 82]. In excising superfluous scaling degrees of freedom, seeking a minimal ontology, we are forced to work with theories which are inherently frictional in nature; consequently, we dedicate this section to developing the geometrical tools required to describe such systems.
@@
does not necessarily define a pre-multicontact structure. It is therefore necessary that a number of additional criteria be met, which we now introduce. To facilitate our analysis, we introduce the notation $\Xi:=\beta^{*}\omega\in\Omega^{d}(\mathcal{S})$. While $\Theta_{L}$ endows $\mathcal{S}$ with the formal geometrical structure of a pre-multicontact manifold, $\Xi$ serves as a reference object. Given a regular distribution $\mathcal{D}\subset T\mathcal{S}$, we introduce the space of $r$-forms which annihilate all sections of $\mathcal{D}$
@@
Sections of $\mathcal{D}^{R}$ are the Reeb vector fields; we denote this space $\mathfrak{R}:=\Gamma(\mathcal{D}^{R})$. In practice, we shall always work in coordinates such that $\Xi=dx^{0}\wedge\,\cdots\,\wedge dx^{d-1}$, from which it follows that $\textrm{ker}\,\Xi$ consists of those vector fields whose vertical components $V^{\mu}\partial_{x^{\mu}}$ are vanishing. The space $\textrm{Ann}^{d}(\textrm{ker}\,\Xi)$ then comprises all $d$-forms $\xi\in\Omega^{d}(\mathcal{S})$, for which $\iota_{X}\xi=0$ for any choice of vector field of the form
@@
In addition to the Reeb distribution, we also have the characteristic distribution $\mathcal{C}$, defined as the intersection $\mathcal{C}:=\textrm{ker}\,\Xi\cap\textrm{ker}\,\Theta_{L}\cap\textrm{ker}\,d\Theta_{L}$. With these geometrical constructions in-hand, the conditions under which the triple $(\mathcal{S},\Theta_{L},\Xi)$ constitutes a pre-multicontact manifold are that, for some $0<k\leqslant n(1+d)$
@@
We have excluded the possibility that $k=0$, as this is precisely the condition for $(\mathcal{S},\Theta_{L},\Xi)$ to be multicontact, and not pre-multicontact. In practice, when working in adapted coordinates, the local Reeb fields are deduced from $\iota_{R_{\mu}}\Theta_{L}=d^{d-1}x_{\mu}$; however, note that this does not determine the $R_{\mu}$ uniquely, for it is possible to add to $R_{\mu}$ an element of the characteristic distribution, that is to say $\mathfrak{R}=\textrm{span}\hskip 1.13809pt(R_{\mu})+\mathcal{C}$.
To any action-dependent Lagrangian, we associate a dissipation form $\sigma_{\Theta_{L}}\in\Omega^{1}(\mathcal{S})$, expressed locally as
@@
$\Omega_{L}:=d\Theta_{L}+\sigma_{\Theta_{L}}\wedge\Theta_{L}=d\Theta_{L}-\frac{\partial L}{\partial s^{\mu}}\,dx^{\mu}\wedge\,\Theta_{L}$ (80)